# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 4.5. Relationship Between Fixed-Point and Gradient Approaches
Let $(\mathbf{X}^{s},\mathrm{p}^{s})$, for $s\in[S]$, and $(\mathbf{Y},\mathrm{q})$ be representatives of the classes $\{[\mathbb{X}^{s}]\}_{s\in[S]}$ (templates) and $[\mathbb{Y}]$ (input shape) in $\mathcal{GM}$. Let us preprocess them obtaining $(\mathbf{X}^{s}_{b},\mathrm{q}_{b})$, $s\in[S]$, and $(\mathbf{Y}_{b},\mathrm{q}_{b})$ in $\mathbb{R}^{M_{b}\times M_{b}}\times\mathcal{P}_{M_{b}}$, using the blow-up Algorithm 3. Consider GW OT plans $\pi^{s}_{b}=\mathrm{diag}(\mathrm{q}_{b})$ for each problem $GW((\mathbf{X}^{s}_{b},\mathrm{q}_{b}),(\mathbf{Y}_{b},\mathrm{q}_{b}))$, with $s\in[S]$. Finally, in the approach outlined in Section 3, consider $M_{b}$ and $\mathrm{q}_{b}$ in place of $M$ and $\mathrm{q}$. We observe that expression (13) becomes:
$\frac{1}{\mathrm{q}_{b}(\mathrm{q}_{b})^{T}}\odot\sum_{s\in[S]}\lambda_{s}(\pi^{s}_{b})^{T}\,\mathbf{X}^{s}_{b}\,\pi^{s}_{b}=\frac{1}{\mathrm{q}_{b}(\mathrm{q}_{b})^{T}}\odot\sum_{s\in[S]}\lambda_{s}\mathrm{diag}(\mathrm{q}_{b})\,\mathbf{X}^{s}_{b}\,\mathrm{diag}(\mathrm{q}_{b})=\sum_{s\in[S]}\lambda_{s}\mathbf{X}^{s}_{b}.$ (38)
@@
Moreover, when formulating the analysis problem (16) (or, equivalently, (19) under the GW-BCM assumption), we have the flexibility to consider any divergence. Therefore, if instead of considering the classical Frobenius norm $d_{a}(A,B)=\|A-B\|_{\mathrm{Frob}}$ in Theorem 3.6, we had used the variant of the Frobenius norm with weight $\mathrm{q}_{b}\otimes\mathrm{q}_{b}$, we would have obtained that the convex quadratic problems (21) and (35) coincide, as $\mathcal{Q}=\mathcal{A}$ (to compare their expressions given in (21) and (34), using that, from (38), one has $F(\mathbf{Y}_{b},s)=\mathbf{X}_{b}^{s}$). Indeed, as functions from $\mathbb{R}^{S}$ to $[0,\infty)$, we would have $\Theta(\lambda)=\|G_{\lambda}([\mathbb{Y}])\|^{2}_{\mathrm{Tan}_{[\mathbb{Y}]}}$ (compare (20) and 4.6). In particular, the linear optimization problems (22) and (36) would coincide; that is, $K=L$ and $b=c$, as defined in (23) and (37). If one tries to reformulate problem (19) to coincide exactly with Section 4.4, one needs to compute $\mathrm{q}_{b}$ in advance. This involves solving one GW problem per template (as part of the blow-up procedure). The blow-up distribution $\mathrm{q}_{b}$ depends on $(\mathbf{Y},\mathrm{q})$ but but generally differs from $\mathrm{q}$.
@@ 5. Experimental Results
In this section, we experimentally assess the performance of our algorithms for solving the GW analysis problem. We also present applications to classification. Additional use cases involving point cloud corruptions are provided in App. D.4.
@@ 5.1. Reconstruction Performance of Algorithms 2 and 4 Under the GW-BCM
In this section, we demonstrate the performance of the proposed GW analysis algorithms by applying them to 3D point cloud data: Figures 6 and 7 illustrate the application of Algorithms 2 and 4 on the ModelNet40 - Princeton 3D Object Dataset [70] consisting of 3D CAD models of 40 common object categories such as chairs, tables, airplanes, lamps, and more. In App. D.1 we provide additional 2D examples (Figures 15 and 16).
From the object category ‘airplane’ we select $S=3$ templates and sample them to generate 3D point clouds with (varying) numbers of points $N^{s}$, $s=1,2,3$ (see Figure 11 in App. D). We compute the associated $N^{s}\times N^{s}$ pairwise Euclidean distance matrices $\mathbf{X}^{s}$ within each point cloud ($s=1,2,3$). We assign to each point uniform mass, that is, $\mathrm{p}^{s}=\frac{1}{N^{s}}(1,\dots,1)\in\mathbb{R}^{N^{s}}$, for $s=1,2,3$. We consider either the uniform vector $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})\in\Delta_{2}$ or a random vector $\lambda\in\Delta_{2}$ drawn from a Dirichlet distribution with uniform parameters $\alpha=(1,1,1)$, which corresponds to the uniform distribution over the 2-simplex. For such a $\lambda$ and the selected templates, we synthesize a GW barycenter $M\times M$ matrix $\mathbf{Y}$ by using the function ot.gromov.gromov_barycenters from the POT Library [21] with a predetermined size $M$ and uniform probability $\mathrm{q}=\frac{1}{M}(1,\dots,1)\in\mathbb{R}^{M}$, which is based on a fixed-point algorithm in the spirit of Algorithm 1. Then, we apply Algorithm 2 or Algorithm 4 to the synthetic GW barycenter, $\mathrm{q}$, and the templates $\{(\mathbf{X}^{1},\mathrm{p}^{1}),(\mathbf{X}^{2},\mathrm{p}^{2}),(\mathbf{X}^{3},\mathrm{p}^{3})\}$ to get the estimate $\widetilde{\lambda}\approx\lambda$.
In Figure 6 we consider $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, and after estimation, we reconstruct the shape as:
@@
    $\mathbf{Y}_{\mathrm{bary}}:=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot\sum_{s=1}^{S}\widetilde{\lambda}_{s}(\pi(\mathbf{Y},s))^{T}\mathbf{X}^{s}\pi(\mathbf{Y},s)\in\mathbb{R}^{M\times M},$ (39)
@@
    $\mathbf{Y}_{\mathrm{barybu}}:=\sum_{s=1}^{S}\widetilde{\lambda}_{s}\mathbf{X}^{s}_{b}\in\mathbb{R}^{M_{b}\times M_{b}},$ (40)
@@
For visualization, we apply the classical multidimesional scaling embedding (MDS) [37] with 3 components to the synthetic GW barycenter matrix $\mathbf{Y}$ and the recovered matrices $\mathbf{Y}_{\mathrm{bary}}$ and $\mathbf{Y}_{\mathrm{barybu}}$ given by the two approaches (39) and (40), respectively. For assessing the reconstruction accuracy, we use the function ot.gromov.gromov_wasserstein from the POT Library to compute $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\mathrm{bary}},\mathrm{q}))$ and $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\mathrm{barybu}},\mathrm{q}_{b}))$.
In Figure 7 we estimate GW barycentric coordinates $\widetilde{\lambda}\approx\lambda$ for synthetic GW barycenters built with randomly generated $\lambda$. The only difference between Figures 7a and 7b is that, in the former, the point cloud templates are uniformly sampled, whereas in the latter, the sampling rates vary across templates. Additional visualizations are Figures 12, 13, 14.
Figure 6. Reconstructions of synthetic GW barycenter. Top Left: MDS of synthetic GW barycenter with 3 templates and $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, computed via ot.gromov.gromov_barycenters (POT Library) with $M=400$ points and uniform mass $\mathrm{q}$. Bottom Left: MDS of the matrix given by (39) after applying Algorithm 2 to estimate $\widetilde{\lambda}\approx\lambda$ (error of order $10^{-14}$). GW distance between shapes in top and bottom left panels: $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\textrm{bary}},\mathrm{q}))\approx 10^{-16}$. Top Right: MDS of the synthetic GW barycenter after blow-up (i.e., using Algorithm 3 with synthesized GW barycenter (top left) and templates). GW distance between the two shapes in the top panel: $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{b},\mathrm{q}_{b}))\approx 10^{-16}$. Bottom Right: MDS of the matrix given by (40) after applying Algorithm 4 to estimate $\widetilde{\lambda}\approx\lambda$ (error of order $10^{-3}$). GW distance between shapes in top left and bottom right panels: $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\textrm{barybu}},\mathrm{q}_{b}))\approx 10^{-5}$.
a We use 3 templates, each uniformly sampled at 150 points. GW barycenters with $M=150$ points were synthesized using ot.gromov.gromov_barycenters from the POT Library.
b We use the 3 templates, sampled at varying rates ranging from 50 to 150 points. GW barycenters with $M=100$ points were synthesized using ot.gromov.gromov_barycenters from the POT Library.
Figure 7. Depiction of 10 different randomly generated $\lambda\in\Delta_{2}$ (blue dots) and their estimation (red crosses) from synthesized GW barycenters using Algorithm 2 (left) and Algorithm 4 (right). Each $\lambda$ is represented as a convex combination of the vertices of an equilateral triangle.
5.2. Classification
Consider the task of identifying a sample’s class label using small number of labeled objects (templates) from the dataset. 3 provides a visualization of how GW barycentric coordinates capture meaningful structure in the data: they separate (or cluster) the digit classes 0 and 1 in the Point Cloud MNIST 2D dataset [23]. The slight observed overlap suggests the presence of shared geometric features among certain point clouds from the two classes. For classification, after estimating the GW barycentric coordinates $\lambda$ of each sample, the predicted label can be assigned by selecting the index corresponding to the highest coordinate in the vector $\lambda$. Additional visualization is in Figure 21.
We now present a similar technique using t-SNE. Additional methods are given in App. D.3, all leveraging GW barycentric coordinates via Algorithms 2 and 4.
5.2.1. t-SNE in the GW Barycentric Coordinate Space – Experimental Setup
Assume there are $C$ classes in a dataset, and consider $k$ reference objects from each class.
  1. (1)
    Randomly select templates from each class of the dataset ($k$ samples per class).
  2. (2)
    For each sample in the dataset, compute GW barycentric coordinates with respect to the selected templates by applying either Algorithm 2 or Algorithm 4. Each sample is thus represented by a coordinate vector $\lambda\in\mathbb{R}^{S}$, where $S=k\times C$.
  3. (3)
    Apply t-SNE (t-distributed Stochastic Neighbor Embedding) as a dimensionality reduction from $\mathbb{R}^{S}$ to $\mathbb{R}^{2}$, and visualize the result, coloring points by class label.
  4. (4)
    Perform clustering using using K-Means (with $C$ number of classes) in the embedded space provided by t-SNE. Compute accuracy by comparing the predicted labels from clustering with the true labels. To assess whether the clustering correctly groups similar digits and to highlight common misclassification patterns, compute the confusion matrix.
As a proof-of-concept, we apply this methodology to the Point Cloud MNIST 2D dataset [23], considering $C=2$ classes (digits 0 and 4) and $k=1,3,6$ template objects from each class. Figures 8 and 9 show the resulting t-SNE embeddings in two dimensions (for 1800 random samples), while the associated K-Means clustering and confusion matrices are provided in App. D (Figures 17, 18, 19, and 20). Algorithms 2 and 4 perform comparably; however, Algorithm 4 is slower due to the preprocessing step (Algorithm 3) required to align samples with the templates. As more templates per class are included, the separation between classes in the t-SNE embedding becomes more pronounced.

### v2

@@ 4.5. Relationship Between Fixed-Point and Gradient Approaches
Let $(\mathbf{X}^{s},\mathrm{p}^{s})$, for $s\in[S]$, and $(\mathbf{Y},\mathrm{q})$ be representatives of the classes $\{[\mathbb{X}^{s}]\}_{s\in[S]}$ (templates) and $\mathbb{Y}]$ (input shape) in $\mathcal{GM}$. Let us preprocess them obtaining $(\mathbf{X}^{s}_{b},\mathrm{q}_{b})$, $s\in[S]$, and $(\mathbf{Y}_{b},\mathrm{q}_{b})$ in $\mathbb{R}^{M_{b}\times M_{b}}\times\mathcal{P}_{M_{b}}$, using the blow-up Algorithm 3. Consider GW OT plans $\pi^{s}_{b}=\mathrm{diag}(\mathrm{q}_{b})$ for each problem $GW((\mathbf{X}^{s}_{b},\mathrm{q}_{b}),(\mathbf{Y}_{b},\mathrm{q}_{b}))$, with $s\in[S]$. Finally, in the approach outlined in Section 3, consider $M_{b}$ and $\mathrm{q}_{b}$ in place of $M$ and $\mathrm{q}$. We observe that expression (10) becomes:
(37) $\frac{1}{\mathrm{q}_{b}(\mathrm{q}_{b})^{T}}\odot\sum_{s\in[S]}\lambda_{s}(\pi^{s}_{b})^{T}\,\mathbf{X}^{s}_{b}\,\pi^{s}_{b}=\frac{1}{\mathrm{q}_{b}(\mathrm{q}_{b})^{T}}\odot\sum_{s\in[S]}\lambda_{s}\mathrm{diag}(\mathrm{q}_{b})\,\mathbf{X}^{s}_{b}\,\mathrm{diag}(\mathrm{q}_{b})=\sum_{s\in[S]}\lambda_{s}\mathbf{X}^{s}_{b}.$
@@
Moreover, when formulating the analysis problem (13) (or, equivalently, (16) under the GW-BCM assumption), we have the flexibility to consider any divergence. Therefore, if instead of considering the classical Frobenius norm $d_{a}(A,B)=\|A-B\|_{\mathrm{Frob}}$ in Theorem 3.6, we had used the variant of the Frobenius norm with weight $\mathrm{q}_{b}\otimes\mathrm{q}_{b}$, we would have obtained that the convex quadratic problems (18) and (34) coincide, as $\mathcal{Q}=\mathcal{A}$ (to compare their expressions given in (18) and (33), using that, from (37), one has $F(\mathbf{Y}_{b},s)=\mathbf{X}_{b}^{s}$). Indeed, as functions from $\mathbb{R}^{S}$ to $0,\infty)$, we would have $\Theta(\lambda)=\|G_{\lambda}([\mathbb{Y}])\|^{2}_{\mathrm{Tan}_{[\mathbb{Y}]}}$ (compare (17) and Proposition 4.6). In particular, the linear optimization problems (19) and (35) would coincide; that is, $K=L$ and $b=c$, as defined in (20) and (36). If one tries to reformulate problem (16) to coincide exactly with Section 4.4, one needs to compute $\mathrm{q}_{b}$ in advance. This involves solving one GW problem per template (as part of the blow-up procedure). The blow-up distribution $\mathrm{q}_{b}$ depends on $(\mathbf{Y},\mathrm{q})$ but but generally differs from $\mathrm{q}$.
Remark 4.9 (Compatibility Test for GW Synthesis Algorithms).
Theorem 4.7 provides a necessary condition for being a “true” GW barycenter. Precisely, if $\mathbb{Y}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$ with corresponding coordinates $\lambda_{\mathbb{Y}}\in\Delta_{S-1}$, then
(38) $\|\nabla G_{\lambda}([\mathbb{Y}])\|_{\mathrm{Tan}_{[\mathbb{Y}]}}^{2}=\lambda_{\mathbb{Y}}^{T}\,\mathcal{A}\,\lambda_{\mathbb{Y}}=0,$
where the $S\times S$ matrix $\mathcal{A}$ depends on the fixed templates $\{\mathbb{X}^{s}\}_{s=1}^{S}$ and on the input $\mathbb{Y}$, and requires preforming the blow-up representation (see (33)). This test, however, does not provide a sufficient guarantee for being a point synthesized through (2). By definition, it is instead necessary and sufficient for addressing if a network belongs to $\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$. In particular, when generating synthetic barycenters (e.g., via Algorithm 1), one can test through (38) whether the obtained object is, at least, a weak GW barycenter.
@@ 5. Experimental Results
In this section, we experimentally assess the performance of our algorithms for solving the GW analysis problem. We also present applications to classification. Additional use cases involving point cloud corruptions are provided in Suppl. Mat. D.2.
@@ 5.1. Reconstruction Performance of Algorithms 2 and 4 Under the GW-BCM
In this section, we demonstrate the performance of the proposed GW analysis algorithms by applying them to 3D point cloud data: Figures 6 and 7 illustrate the application of Algorithms 2 and 4 on the ModelNet40 - Princeton 3D Object Dataset [74] consisting of 3D CAD models of 40 common object categories such as chairs, tables, airplanes, lamps, and more.
From the object category ‘airplane’ we select $S=3$ templates and sample them to generate 3D point clouds with (varying) numbers of points $N^{s}$, $s=1,2,3$ (Fig. 6). We compute the associated $N^{s}\times N^{s}$ pairwise Euclidean distance matrices $\mathbf{X}^{s}$ within each point cloud ($s=1,2,3$). We assign to each point uniform mass, that is, $\mathrm{p}^{s}=\frac{1}{N^{s}}(1,\dots,1)\in\mathbb{R}^{N^{s}}$, for $s=1,2,3$. We consider either the uniform vector $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})\in\Delta_{2}$ or a random vector $\lambda\in\Delta_{2}$ drawn from a Dirichlet distribution with uniform parameters $\alpha=(1,1,1)$, which corresponds to the uniform distribution over the 2-simplex. For such a $\lambda$ and the selected templates, we synthesize a GW barycenter $M\times M$ matrix $\mathbf{Y}$ by using the function ot.gromov.gromov_barycenters from the POT Library [22] with a predetermined size $M$ and uniform probability $\mathrm{q}=\frac{1}{M}(1,\dots,1)\in\mathbb{R}^{M}$, which is based on a fixed-point algorithm in the spirit of Algorithm 1. Then, we apply Algorithm 2 or Algorithm 4 to the synthetic GW barycenter, $\mathrm{q}$, and the templates $\{(\mathbf{X}^{1},\mathrm{p}^{1}),(\mathbf{X}^{2},\mathrm{p}^{2}),(\mathbf{X}^{3},\mathrm{p}^{3})\}$ to get the estimate $\widetilde{\lambda}\approx\lambda$.
In Fig. 6 we consider $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, and after estimation, we reconstruct the shape as:
@@
    (39) $\mathbf{Y}_{\mathrm{bary}}:=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot\sum_{s\in[S]}\widetilde{\lambda}_{s}(\pi(\mathbf{Y},s))^{T}\mathbf{X}^{s}\pi(\mathbf{Y},s)\in\mathbb{R}^{M\times M},$
@@
    (40) $\mathbf{Y}_{\mathrm{barybu}}:=\sum_{s\in[S]}\widetilde{\lambda}_{s}\,\mathbf{X}^{s}_{b}\in\mathbb{R}^{M_{b}\times M_{b}},$
@@
For visualization, we apply the classical multidimensional scaling embedding (MDS) [39] with 3 components to the synthetic GW barycenter matrix $\mathbf{Y}$ and the recovered matrices $\mathbf{Y}_{\mathrm{bary}}$ and $\mathbf{Y}_{\mathrm{barybu}}$ given by the two approaches (39) and (40), respectively. For assessing the reconstruction accuracy, we use the function ot.gromov.gromov_wasserstein from the POT Library to compute $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\mathrm{bary}},\mathrm{q}))$ and $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}_{\mathrm{barybu}},\mathrm{q}_{b}))$.
In Fig. 7 we estimate GW barycentric coordinates $\widetilde{\lambda}\approx\lambda$ for synthetic GW barycenters built with randomly generated $\lambda$. We synthesize with input size $M=30$ and notice that when using templates having the same number of nodes (equal to $30$), the blow-up average size (calculated when applying Algorithm 4) is $M_{b}=30$; and when using templates having different number of nodes (between $25$ and $35$), the blow-up average size obtained is $M_{b}\approx 107$ (same order as $4\times 30$, cf. Remark 4.8).
Reconstructions of synthetic GW barycenters using MDS on distance matrices.
$\mathrm{GW}(\text{{(a)}},\text{{(b)}})\approx 10^{-16}$
$\mathrm{GW}(\text{{(a)}},\text{{(c)}})\approx 10^{-16}$
$\mathrm{GW}(\text{{(a)}},\text{{(d)}})\approx 10^{-5}$
Order of GW Errors.
Templates.
$\mathrm{GW}(\text{{(i)}},\text{{(ii)}})\approx 10^{-16}$
$\mathrm{GW}(\text{{(i)}},\text{{(iii)}})\approx 10^{-17}$
$\mathrm{GW}(\text{{(i)}},\text{{(iv)}})\approx 10^{-17}$
Order of GW Errors.
Figure 6. (a) GW barycenter $(\mathbf{Y},\mathrm{q})$ consisting of $M=400$ points and uniform mass $\mathrm{q}$, synthesized with the templates above sampled with different number of points between $300$ and $500$. (b) MDS of the matrix given by (39) (i.e., $(\mathbf{Y}_{\textrm{bary}},\mathrm{q})$) after applying the fixed-point analysis Algorithm 2 to (a) to estimate $\widetilde{\lambda}\approx\lambda$ (error of order $10^{-14}$). (c) The resulting blow-up of (a) (i.e., $(\mathbf{Y}_{b},\mathrm{q}_{b})$) when Algorithm 3 is applied to the templates and (a). The size $M_{b}$ was about $1500$, which is of the same order as $400+3N$ where $300\leq N\leq 500$ (cf. Remark 4.8). (d) MDS of the matrix given by (40) (i.e., $(\mathbf{Y}_{\textrm{barybu}},\mathrm{q}_{b})$) after applying the blow-up analysis Algorithm 4 to (a) to recover $\widetilde{\lambda}\approx\lambda$ (error of order $10^{-3}$). (i)–(iv) Analogous to (a)–(d) but uniformly sampling the templates with $M=400$ points. Errors in $\lambda$ estimation for (ii) and (iv) are of order $10^{-15}$ and $10^{-16}$, respectively. Blow-up size $M_{b}=400$. Difference: (a) does not satisfy the test from Remark 4.9, while (i) passes the test.
Figure 7. Depiction of 100 different randomly generated $\lambda\in\Delta_{2}$ (dots) and their estimation $\widetilde{\lambda}$ (red crosses) from synthesized GW barycenters, using Algorithm 2 (left) and Algorithm 4 (right). Each $\lambda$ is represented as a convex combination of the vertices of an equilateral triangle. We compute mean and standard deviation of the reconstruction error $\lVert\lambda-\widetilde{\lambda}\rVert_{2}$ and of the average wall-clock time (in seconds) of our GW-analysis methods over $100$ independent runs. For each experiment, we use 3 templates, uniformly sampled at 30 points, with uniform distribution of mass across nodes. The 100 GW barycenters, with $M=30$ points and $\mathrm{q}=\frac{1}{30}(1,\dots,1)\in\mathbb{R}^{30}$ each, were synthesized using ot.gromov.gromov_barycenters from the POT Library. Blow-up average size when using Algorithm 2: $M_{b}=30$. Blue dots (85) represent synthesized output that pass the compatibility test in Remark 4.9, whereas green dots do not (15). In the right-hand boxed panel we run 10 experiments with templates sampled at varying rates noticing all the synthetic outputs fail to be critical points of the GW synthesis functional (all green dots).
5.2. Clustering, Classification and Visualization
In this section, we expand on the motivating visualization in Fig. 2, illustrating how GW barycentric coordinates capture meaningful structure in the data: they separate classes while showing slight overlaps when samples shared geometric features. Generally, suppose a dataset contains $C$ classes, and $u$ templates are randomly selected from each class. For each sample, one can compute GW barycentric coordinates with respect to the templates using either Algorithm 2 or Algorithm 4. Each sample is thus represented by a coordinate vector $\lambda\in\mathbb{R}^{S}$, where $S=u\times C$. These coordinates can be used directly for supervised classification (e.g., k-NN) or for unsupervised analysis such as K-Means clustering. To visualize the coordinates when $S>3$, one can apply another dimensionality reduction such as t-SNE to map $\mathbb{R}^{S}$ to $\mathbb{R}^{2}$.
As a proof of concept, we apply this framework to the Point Cloud MNIST 2D dataset [24]. We first consider classification experiments with $C=2$ and $u=1$; results are illustrated in Fig. 8 and reported in Table 1. Algorithms 2 and 4 perform comparably in this setting, although Algorithm 4 incurs additional computational cost due to the preprocessing step (Algorithm 3). To generate Table 1, we treat the $\lambda$-vectors as features and apply a k-NN classifier. We compare against two baselines. First, we consider a naive 1-NN classifier based directly on the GW distance: by using only the selected templates as the training set, each sample $\mathbb{Y}$ is assigned the label of the template $\mathbb{X}^{s}$ minimizing $\mathrm{GW}(\mathbb{X}^{s},\mathbb{Y})$. Second, we consider the GW dictionary-learning approach of [69] (without regularization), implemented in the POT Library [22]. Dictionary atoms (of fixed size $20$, learned through 5 epochs) are learned using the full dataset, and each sample is represented by so-called unmixing weights obtained by solving a GW linear-like approximation problem. This objective differs from the GW barycenter problem and should be viewed as a computationally tractable relaxation. Additional classification experiments can be found in Suppl. Mat. D.1.2.
a Fixed-Point approach – Digits 0&40\&4.
b Blow-Up approach – Digits 0&40\&4.
c Digits 2&52\&5.
Figure 8. (a)-(b): We considered the classes of digits 0 and 4 from the Point Cloud MNIST 2D dataset [24], and selected 900 samples and one template per class. The GW coordinates $(\lambda_{1},\lambda_{2})\in\Delta_{1}$ of each data point are computed using Algorithm 2 in (a) and Algorithm 4 in (b). Since the representation is intrinsically one-dimensional, the plots visualize the empirical distributions of the GW barycentric coordinate $\lambda:=\lambda_{1}\in[0,1]$. Regions of overlap correspond to geometrically ambiguous samples and explain the observed classification errors of k-NN (see Table 1). The mass across nodes in each point cloud is not uniformly distributed, and each point cloud in the dataset has a different number of nodes, varying from 50 to 200 points with an average of 120. The blow-up average size $M_{b}$ from the experiments on the right is $415.5$ (which is of the same order as $3\times 120$, cf. remark 4.8). Subplot (c), where 200 samples of the digits 2 and 5 where considered, indicates that the framework captures meaningful structures in the data since those digits share their internal structure and overlap is expected.
$\lambda$-GW (FP) + k-NN $\lambda$-GW (BU) + k-NN GW 1-NN GW dict + k-NN
Accuracy (mean $\pm$ std) 0.94 $\pm$ 0.009 0.95 $\pm$ 0.012 0.89 $\pm$ 0.016 0.93 $\pm$ 0.016
Time (s) 92.67 ($\lambda$-GW coord) 147.23 ($\lambda$-GW coord) 97.17 (avg run) 50.21 (learn atoms) + 10.05 (unmix weights)
Table 1. Classification performance and runtime comparison on the Point Cloud MNIST 2D dataset [24] for digits 0 and 4 with $200$ samples each. For columns 2, 3, and 5, each sample is represented either by GW barycentric coordinates $\lambda$ computed via the fixed-point (FP) approach (Algorithm 2), the blow-up (BU) approach (Algorithm 4), or by GW dictionary unmixing weights [69], respectively, and then classified using k-NN in the resulting feature space. For k-NN methods, we use repeated stratified 70/30 train/test splits (20 outer repetitions) and select $k$ by 5-fold stratified cross-validation (CV) performed on the training split within each outer repetition. Specifically, $k$ is chosen across $\{1,3,5,7,9\}$ to maximize the mean CV accuracy; the classifier is then retrained on the training set and evaluated on the test set from the corresponding outer split. In column 4 (GW 1-NN), each sample is assigned the label of the template that minimizes the GW distance.
We then consider unsupervised analysis and embedding visualization using multiple templates per class. For $u=3,6$, we perform K-Means on the GW coordinate space and apply t-SNE for visualization; the resulting embeddings are shown in Fig. 9. As illustrated in Figures 9a and 9b, increasing the number of templates per class leads to clearer separation between classes in the embedded space. Additional visualizations with three classes ($C=3$) are provided in Fig. 10 and Fig. 13 in Suppl. Mat. (with $u=1$).
a 3 templates per class–FP.
b 6 templates per class–FP.
c 6 templates per class–BU.
Figure 9. Visualization through t-SNE of estimated $\lambda$ coordinates via Algorithm 2 (a)-(b) and Algorithm 4 (c). For (a) and (b) we consider 1800 total samples, and 500 for (c). Multiple templates per class: (a) 3; (b) 6; (c) 6. Clustering performance via K-Means in $\lambda$-space: Adjusted Rand Index for FP analysis, (a) 0.76, (b) 0.81; for BU analysis, 0.86 when using 3 templates per class.